# III. Interacting Dark Energy: Summary of Models, Pathologies, and Constraints

### v1

@@ 1 Introduction
In our two companion papers titled “I. Linear Interacting Dark Energy: Analytical Solutions and Theoretical Pathologies” and “II. Non-Linear Interacting Dark Energy: Analytical Solutions and Theoretical Pathologies,” we provided an in-depth study of the background dynamics of eight different interacting dark energy (IDE) models. The eight interaction kernels studied are phenomenological and were chosen due to their simplicity, which allowed new analytical solutions for the evolution of dark matter (DM) and dark energy (DE) to be obtained from the modified conservation equations. These specific interaction kernels are also some of the most widely studied (see our companion papers and the references therein), which prompted us to fill in gaps within the literature, especially with regard to the often-overlooked presence of both negative energy densities and future big rip singularities. We also discussed how these models address the coincidence problem, and the possibility of phantom crossing appearing in either the effective DE equation of state or the reconstructed dynamical DE equation of state, such that at some point we have either $w^{\rm eff}_{\rm de}(z)=-1$ or $\tilde{w}(z)=-1$.
@@
    The Hubble tension: The $4\sigma-6\sigma$ discrepancy in the estimation of the present expansion rate $H_{0}$ from late-time probes such as Type Ia Supernova and early-time probes such as CMB. The potential of IDE models to address this tension has caused a resurgence in their popularity in recent years [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60].
@@ 2.1.1 Linear IDE model 1: $Q=3H(\delta_{\text{dm}}\rho_{\text{dm}}+\delta_{\text{de}}\rho_{\text{de}})$
This is the most general IDE model we studied, with the four following linear models being special cases of this interaction kernel, depending on the sign and magnitude of coupling constants $\delta_{\text{dm}}$ and $\delta_{\text{de}}$. Due to the presence of a determinant $\Delta$ in both $\rho_{\text{dm}}$ and $\rho_{\text{de}}$, care should be taken to avoid imaginary values, using the conditions provided in Table 1. Negative energies may also be avoided using the set of three conditions given in Table 2. The densities of DM and DE evolve with the scale factor according to the expressions (2.2) and (2.3), where $w$ is the DE equation of state. This interaction has been studied in [120, 121, 122, 14, 9, 123, 11, 124, 125, 126, 127, 124, 128, 129, 12, 130, 131, 132, 8, 133, 134, 135, 136, 137, 138, 139].
@@
$\begin{split}h(z)=\Bigg{\{}-&\frac{1}{2\Delta}\Bigl{[}\Omega_{\text{(de,0)}}\Bigl{(}\delta_{\text{dm}}-\delta_{\text{de}}+w-\Delta\Bigr{)}+\Omega_{\text{(dm,0)}}\Bigl{(}\delta_{\text{dm}}-\delta_{\text{de}}-w-\Delta\Bigr{)}\Bigr{]}(1+z)^{-\tfrac{3}{2}\Bigl{(}\delta_{\text{dm}}-\delta_{\text{de}}-w-2+\Delta\Bigr{)}}\\[2.84526pt] +&\frac{1}{2\Delta}\Bigl{[}\Omega_{\text{(de,0)}}\Bigl{(}\delta_{\text{dm}}-\delta_{\text{de}}+w+\Delta\Bigr{)}+\Omega_{\text{(dm,0)}}\Bigl{(}\delta_{\text{dm}}-\delta_{\text{de}}-w+\Delta\Bigr{)}\Bigr{]}(1+z)^{-\tfrac{3}{2}\Bigl{(}\delta_{\text{dm}}-\delta_{\text{de}}-w-2-\Delta\Bigr{)}}\\[2.84526pt] +&\Omega_{\text{bm,0}}(1+z)^{3}+\Omega_{\text{r,0}}(1+z)^{4}\Bigg{\}}^{\tfrac{1}{2}}.\end{split}$ (2.5)
@@ 2.1.2 Linear IDE model 2: $Q=3H\delta(\rho_{\text{dm}}+\rho_{\text{de}})$
This interaction changes the dynamics in both the distant past and future, during DM and DE domination, respectively. Since $Q\neq 0$ when either $\rho_{\text{dm}}$ or $\rho_{\text{de}}$ becomes zero, there is no mechanism to avoid negative energies for this interaction. This model has both negative DM and DE densities in the iDMDE regime, but this can be avoided with a sufficiently small interaction in the iDEDM regime, as given by the conditions in Table 2. Due to the presence of a determinant $\Delta$ in both $\rho_{\text{dm}}$ and $\rho_{\text{de}}$, imaginary values may appear, but these can be avoided using the conditions provided in Table 1. The densities of DM and DE evolve with the scale factor according to the expressions (2.6) and (2.7). This interaction was previously studied in [140, 141, 142, 125, 143, 7, 142, 128, 129, 12, 10, 144, 134, 145, 146, 147, 148, 130, 132, 149, 150, 151, 152, 153, 154, 155, 156, 141, 157, 158, 159, 151, 141, 160].
@@
$\begin{split}h(z)=\Bigg{\{}-&\frac{1}{2\Delta}\Bigl{[}\Omega_{\text{(de,0)}}(w-\Delta)+\Omega_{\text{(dm,0)}}(-w-\Delta)\Bigr{]}(1+z)^{-\tfrac{3}{2}(-w-2+\Delta)}\\[2.84526pt] +&\frac{1}{2\Delta}\Bigl{[}\Omega_{\text{(de,0)}}(w+\Delta)+\Omega_{\text{(dm,0)}}(-w+\Delta)\Bigr{]}(1+z)^{-\tfrac{3}{2}(-w-2-\Delta)}\\[2.84526pt] +&\Omega_{\text{bm,0}}(1+z)^{3}+\Omega_{\text{r,0}}(1+z)^{4}\Bigg{\}}^{\tfrac{1}{2}}.\end{split}$ (2.9)
@@ 2.1.3 Linear IDE model 3: $Q=3H\delta(\rho_{\text{dm}}-\rho_{\text{de}})$
$\begin{split}\rho_{\text{de}}=&+\frac{2\delta-w+\Delta}{4w\Delta}\Bigl{[}\rho_{\text{(de,0)}}(2\delta+w-\Delta)+\rho_{\text{(dm,0)}}(2\delta-w-\Delta)\Bigr{]}a^{\tfrac{3}{2}(2\delta-w-2+\Delta)}\\[2.84526pt] -&\frac{2\delta-w-\Delta}{4w\Delta}\Bigl{[}\rho_{\text{(de,0)}}(2\delta+w+\Delta)+\rho_{\text{(dm,0)}}(2\delta-w+\Delta)\Bigr{]}a^{\tfrac{3}{2}(2\delta-w-2-\Delta)},\end{split}$ (2.11)
@@
$\begin{split}h(z)=\Bigg{\{}-&\frac{1}{2\Delta}\Bigl{[}\Omega_{\text{(de,0)}}(2\delta+w-\Delta)+\Omega_{\text{(dm,0)}}(2\delta-w-\Delta)\Bigr{]}(1+z)^{-\tfrac{3}{2}(2\delta-w-2+\Delta)}\\[2.84526pt] +&\frac{1}{2\Delta}\Bigl{[}\Omega_{\text{(de,0)}}(2\delta+w+\Delta)+\Omega_{\text{(dm,0)}}(2\delta-w+\Delta)\Bigr{]}(1+z)^{-\tfrac{3}{2}(2\delta-w-2-\Delta)}\\[2.84526pt] +&\Omega_{\text{bm,0}}(1+z)^{3}+\Omega_{\text{r,0}}(1+z)^{4}\Bigg{\}}^{\tfrac{1}{2}}.\end{split}$ (2.13)
@@ 2.1.4 Linear IDE model 4: $Q=3H\delta\rho_{\text{dm}}$
This interaction mostly changes the dynamics in the distant past during DM domination. For this interaction, $Q=0$ if $\rho_{\text{dm}}=0$; therefore, we can guarantee that $\rho_{\text{dm}}\geq 0$ at all times. In contrast, this interaction will have negative DE densities in the iDMDE regime, but this can be avoided with a sufficiently small interaction in the iDEDM regime, as given by the conditions in Table 2. The densities of DM and DE evolve with the scale factor according to the expressions (2.14) and (2.15). This interaction was previously studied in [6, 14, 165, 142, 123, 166, 167, 125, 143, 36, 168, 126, 36, 169, 142, 128, 129, 170, 12, 171, 172, 133, 173, 174, 51, 52, 54, 130, 132, 175, 149, 150, 176, 14, 151, 156, 141, 177, 178, 157, 158, 159, 179, 151, 141, 180, 181, 152, 153, 154, 155, 173, 182, 183, 184, 185, 174, 186, 187, 145, 188, 189, 186, 190, 148, 191, 187, 192, 193, 194, 195, 196].
@@
$\begin{split}h(z)=\Bigg{\{}&\Omega_{\text{(dm,0)}}(1+z)^{3(1-\delta)}\\[2.84526pt] &+\Biggl{(}\Omega_{\text{(de,0)}}+\Omega_{\text{(dm,0)}}\left(\frac{\delta}{\delta+w}\right)\Bigl{[}1-(1+z)^{-3(\delta+w)}\Bigr{]}\Biggr{)}(1+z)^{3(w+1)}\\[2.84526pt] &+\Omega_{\text{bm,0}}(1+z)^{3}+\Omega_{\text{r,0}}(1+z)^{4}\Bigg{\}}^{\tfrac{1}{2}}.\end{split}$ (2.16)
@@ 2.1.5 Linear IDE model 5: $Q=3H\delta\rho_{\text{de}}$
This interaction mostly changes the dynamics for the late-time expansion and the distant future during DE domination. For this interaction, $Q=0$ if $\rho_{\text{de}}=0$; therefore, we can guarantee that $\rho_{\text{de}}\geq 0$ at all times. In contrast, this interaction will have negative DM densities in the iDMDE regime, but this can be avoided with a sufficiently small interaction in the iDEDM regime, as given by the conditions in Table 2. The densities of DM and DE evolve with the scale factor according to the expressions (2.17) and (2.18). This interaction was previously studied in [142, 129, 14, 36, 142, 123, 129, 197, 198, 125, 199, 200, 201, 143, 36, 126, 127, 36, 169, 142, 128, 129, 170, 12, 202, 203, 171, 172, 204, 199, 133, 205, 206, 61, 48, 174, 207, 63, 208, 209, 62, 52, 64, 49, 54, 210, 211, 55, 56, 65, 212, 59, 213, 130, 132, 149, 175, 150, 176, 14, 156, 214, 215, 177, 178, 157, 158, 159, 195, 216, 204, 179, 151, 217, 55, 141, 180, 181, 152, 153, 154, 155, 182, 61, 218, 183, 185, 219, 207, 174, 186, 216, 59, 220, 187, 145, 221, 222, 148, 223, 192, 47, 224, 225, 226, 195, 227, 228, 66].
@@
$\begin{split}h(z)=\Bigg{\{}&\Biggl{(}\Omega_{\text{(dm,0)}}+\Omega_{\text{(de,0)}}\left(\frac{\delta}{\delta+w}\right)\Bigl{[}1-(1+z)^{3(\delta+w)}\Bigr{]}\Biggr{)}(1+z)^{3}\\[2.84526pt] &+\Omega_{\text{(de,0)}}(1+z)^{3(\delta+w+1)}+\Omega_{\text{bm,0}}(1+z)^{3}+\Omega_{\text{r,0}}(1+z)^{4}\Bigg{\}}^{\tfrac{1}{2}}.\end{split}$ (2.19)
@@ 2.2 Non-linear IDE models
$\begin{split}h(z)=\Bigg{\{}&\Bigl{[}\Omega_{\text{(dm,0)}}(1+z)^{3(1-\delta)}+\Omega_{\text{(de,0)}}(1+z)^{3(1+w)}\Bigr{]}\left[\frac{1+\left(\tfrac{\Omega_{\text{(dm,0)}}}{\Omega_{\text{(de,0)}}}\right)(1+z)^{-3(w+\delta)}}{1+\left(\tfrac{\Omega_{\text{(dm,0)}}}{\Omega_{\text{(de,0)}}}\right)}\right]^{-\tfrac{\delta}{w+\delta}}\\[2.84526pt] &+\Omega_{\text{bm,0}}(1+z)^{3}+\Omega_{\text{r,0}}(1+z)^{4}\Bigg{\}}^{\tfrac{1}{2}}.\end{split}$ (2.22)
@@
Similar to the kernel $Q=3H\delta\rho_{\text{dm}}$, this interaction primarily affects the dynamics in the distant past during DM domination but has a smaller impact on the late-time dynamics during DE domination. For this interaction, $Q=0$ if $\rho_{\text{dm}}=0$; therefore, we can guarantee that $\rho_{\text{dm}}\geq 0$ at all times. In contrast, this interaction leads to negative DE densities in the iDMDE regime, which may be avoided with a small interaction in the iDEDM regime, as specified by the conditions in Table 2. The densities of DM and DE evolve with the scale factor according to the expressions (2.23) and (2.24). This interaction was previously studied in [229, 231, 14, 133, 134, 135, 168, 14, 156].
@@
$\begin{split}h(z)=\Bigg{\{}&\Biggl{[}\Omega_{\text{(dm,0)}}+\Omega_{\text{(de,0)}}\left(\frac{\Bigl{[}w+\delta\left(\tfrac{\Omega_{\text{(dm,0)}}}{\Omega_{\text{(de,0)}}}\right)\Bigr{]}(1+z)^{3w}-\delta\left(\tfrac{\Omega_{\text{(dm,0)}}}{\Omega_{\text{(de,0)}}}\right)}{w}\right)\Biggr{]}(1+z)^{3\left(1-\tfrac{w\delta}{w-\delta}\right)}\\[2.84526pt] &\times\left[\frac{\Bigl{[}w+\delta\left(\tfrac{\Omega_{\text{(dm,0)}}}{\Omega_{\text{(de,0)}}}\right)\Bigr{]}(1+z)^{3w}+\left(\tfrac{\Omega_{\text{(dm,0)}}}{\Omega_{\text{(de,0)}}}\right)(w-\delta)}{w\Bigl{[}1+\tfrac{\Omega_{\text{(dm,0)}}}{\Omega_{\text{(de,0)}}}\Bigr{]}}\right]^{\tfrac{\delta}{w-\delta}}\\[2.84526pt] &+\Omega_{\text{bm,0}}(1+z)^{3}+\Omega_{\text{r,0}}(1+z)^{4}\Bigg{\}}^{\tfrac{1}{2}}.\end{split}$ (2.25)
@@
Similar to the kernel $Q=3H\delta\rho_{\text{de}}$, this interaction primarily affects the dynamics of the late-time expansion and the distant future during DE domination, but has a smaller impact on the past dynamics during DM domination. For this interaction, $Q=0$ if $\rho_{\text{de}}=0$; therefore, we can guarantee that $\rho_{\text{de}}\geq 0$ at all times. In contrast, this interaction leads to negative DM densities in the iDMDE regime, but this can be avoided with a sufficiently small interaction in the iDEDM regime, as specified by the conditions in Table 2. The densities of DM and DE evolve with the scale factor according to the expressions (2.26) and (2.27). This interaction was previously studied in [229, 231, 14, 133, 134, 135, 14, 156].
@@
$\begin{split}h(z)=\Bigg{\{}&\Biggl{[}\Omega_{\text{(dm,0)}}\left(\frac{\Bigl{[}w\left(\tfrac{\Omega_{\text{(dm,0)}}}{\Omega_{\text{(de,0)}}}\right)+\delta\Bigr{]}(1+z)^{-3w}-\delta}{w\left(\tfrac{\Omega_{\text{(dm,0)}}}{\Omega_{\text{(de,0)}}}\right)}\right)+\Omega_{\text{(de,0)}}\Biggr{]}(1+z)^{3\left(1+\tfrac{w^{2}}{w-\delta}\right)}\\[2.84526pt] &\times\left[\frac{\Bigl{[}w\left(\tfrac{\Omega_{\text{(dm,0)}}}{\Omega_{\text{(de,0)}}}\right)+\delta\Bigr{]}(1+z)^{-3w}+w-\delta}{w\left(1+\tfrac{\Omega_{\text{(dm,0)}}}{\Omega_{\text{(de,0)}}}\right)}\right]^{\tfrac{\delta}{w-\delta}}\\[2.84526pt] &+\Omega_{\text{bm,0}}(1+z)^{3}+\Omega_{\text{r,0}}(1+z)^{4}\Bigg{\}}^{\tfrac{1}{2}}.\end{split}$ (2.28)

### v3

@@ 1 Introduction
In our two companion papers titled “I. Linear Interacting Dark Energy: Analytical Solutions and Theoretical Pathologies”[1] and “II. Non-Linear Interacting Dark Energy: Analytical Solutions and Theoretical Pathologies” [2], we provided an in-depth study of the background dynamics of eight different interacting dark energy (IDE) models. The eight interaction kernels studied are phenomenological and were chosen due to their simplicity, which allowed new analytical solutions for the evolution of dark matter (DM) and dark energy (DE) to be obtained from the modified conservation equations. These specific interaction kernels are also some of the most widely studied (see our companion papers and the references therein), which prompted us to fill in gaps within the literature, especially with regard to the often-overlooked presence of both negative energy densities and future big rip singularities. We also discussed how these models address the coincidence problem, and the possibility of phantom crossing appearing in either the effective DE equation of state or the reconstructed dynamical DE equation of state, such that at some point we have either $w^{\rm eff}_{\rm de}(z)=-1$ or $\tilde{w}(z)=-1$.
@@
    The Hubble tension: The $4\sigma-6\sigma$ discrepancy in the estimation of the present expansion rate $H_{0}$ from late-time probes such as Type Ia Supernova and early-time probes such as CMB. The potential of IDE models to address this tension has caused a resurgence in their popularity in recent years [17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64].
@@ 2.1.1 Linear IDE model 1: $Q=3H(\delta_{\text{dm}}\rho_{\text{dm}}+\delta_{\text{de}}\rho_{\text{de}})$
This is the most general IDE model we studied, with the four following linear models being special cases of this interaction kernel, depending on the sign and magnitude of coupling constants $\delta_{\text{dm}}$ and $\delta_{\text{de}}$. Due to the presence of a determinant $\Delta$ in both $\rho_{\text{dm}}$ and $\rho_{\text{de}}$, care should be taken to avoid imaginary values, using the conditions provided in Table 1. Negative energies may also be avoided using the set of three conditions given in Table 2. The densities of DM and DE evolve with the scale factor according to the expressions (2.2) and (2.3), where $w$ is the DE equation of state. This interaction has been studied in [126, 127, 128, 16, 11, 129, 13, 130, 131, 132, 133, 134, 135, 14, 136, 137, 138, 10, 139, 140, 141, 142, 143, 144, 145].
@@
$\begin{split}h(z)=\Bigg\{-&\frac{1}{2\Delta}\Bigl[\Omega_{\text{(de,0)}}\Bigl(\delta_{\text{dm}}-\delta_{\text{de}}+w-\Delta\Bigr)+\Omega_{\text{(dm,0)}}\Bigl(\delta_{\text{dm}}-\delta_{\text{de}}-w-\Delta\Bigr)\Bigr](1+z)^{-\tfrac{3}{2}\Bigl(\delta_{\text{dm}}-\delta_{\text{de}}-w-2+\Delta\Bigr)}\\[2.84526pt] +&\frac{1}{2\Delta}\Bigl[\Omega_{\text{(de,0)}}\Bigl(\delta_{\text{dm}}-\delta_{\text{de}}+w+\Delta\Bigr)+\Omega_{\text{(dm,0)}}\Bigl(\delta_{\text{dm}}-\delta_{\text{de}}-w+\Delta\Bigr)\Bigr](1+z)^{-\tfrac{3}{2}\Bigl(\delta_{\text{dm}}-\delta_{\text{de}}-w-2-\Delta\Bigr)}\\[2.84526pt] +&\Omega_{\text{(bm,0)}}(1+z)^{3}+\Omega_{\text{(r,0)}}(1+z)^{4}\Bigg\}^{\tfrac{1}{2}}.\end{split}$ (2.5)
@@ 2.1.2 Linear IDE model 2: $Q=3H\delta(\rho_{\text{dm}}+\rho_{\text{de}})$
This interaction changes the dynamics in both the distant past and future, during DM and DE domination, respectively. Since $Q\neq 0$ when either $\rho_{\text{dm}}$ or $\rho_{\text{de}}$ becomes zero, there is no mechanism to avoid negative energies for this interaction. This model has both negative DM and DE densities in the iDMDE regime, but this can be avoided with a sufficiently small interaction in the iDEDM regime, as given by the conditions in Table 2. Due to the presence of a determinant $\Delta$ in both $\rho_{\text{dm}}$ and $\rho_{\text{de}}$, imaginary values may appear, but these can be avoided using the conditions provided in Table 1. The densities of DM and DE evolve with the scale factor according to the expressions (2.6) and (2.7). This interaction was previously studied in [146, 147, 131, 148, 9, 134, 135, 14, 12, 149, 140, 150, 151, 152, 153, 136, 138, 154, 155, 156, 157, 158, 159, 160, 161, 147, 162, 163, 164, 156, 147, 165, 166].
@@
$\begin{split}h(z)=\Bigg\{-&\frac{1}{2\Delta}\Bigl[\Omega_{\text{(de,0)}}(w-\Delta)+\Omega_{\text{(dm,0)}}(-w-\Delta)\Bigr](1+z)^{-\tfrac{3}{2}(-w-2+\Delta)}\\[2.84526pt] +&\frac{1}{2\Delta}\Bigl[\Omega_{\text{(de,0)}}(w+\Delta)+\Omega_{\text{(dm,0)}}(-w+\Delta)\Bigr](1+z)^{-\tfrac{3}{2}(-w-2-\Delta)}\\[2.84526pt] +&\Omega_{\text{(bm,0)}}(1+z)^{3}+\Omega_{\text{(r,0)}}(1+z)^{4}\Bigg\}^{\tfrac{1}{2}}.\end{split}$ (2.9)
@@ 2.1.3 Linear IDE model 3: $Q=3H\delta(\rho_{\text{dm}}-\rho_{\text{de}})$
$\begin{split}\rho_{\text{de}}=&+\frac{2\delta-w+\Delta}{4w\Delta}\Bigl[\rho_{\text{(de,0)}}(2\delta+w-\Delta)+\rho_{\text{(dm,0)}}(2\delta-w-\Delta)\Bigr]a^{\tfrac{3}{2}(2\delta-w-2+\Delta)}\\[2.84526pt] &-\frac{2\delta-w-\Delta}{4w\Delta}\Bigl[\rho_{\text{(de,0)}}(2\delta+w+\Delta)+\rho_{\text{(dm,0)}}(2\delta-w+\Delta)\Bigr]a^{\tfrac{3}{2}(2\delta-w-2-\Delta)},\end{split}$ (2.11)
@@
$\begin{split}h(z)=\Bigg\{-&\frac{1}{2\Delta}\Bigl[\Omega_{\text{(de,0)}}(2\delta+w-\Delta)+\Omega_{\text{(dm,0)}}(2\delta-w-\Delta)\Bigr](1+z)^{-\tfrac{3}{2}(2\delta-w-2+\Delta)}\\[2.84526pt] +&\frac{1}{2\Delta}\Bigl[\Omega_{\text{(de,0)}}(2\delta+w+\Delta)+\Omega_{\text{(dm,0)}}(2\delta-w+\Delta)\Bigr](1+z)^{-\tfrac{3}{2}(2\delta-w-2-\Delta)}\\[2.84526pt] +&\Omega_{\text{(bm,0)}}(1+z)^{3}+\Omega_{\text{(r,0)}}(1+z)^{4}\Bigg\}^{\tfrac{1}{2}}.\end{split}$ (2.13)
@@ 2.1.4 Linear IDE model 4: $Q=3H\delta\rho_{\text{dm}}$
This interaction mostly changes the dynamics in the distant past during DM domination. For this interaction, $Q=0$ if $\rho_{\text{dm}}=0$; therefore, we can guarantee that $\rho_{\text{dm}}\geq 0$ at all times. In contrast, this interaction will have past negative DE densities in the iDMDE regime, but this can be avoided with a sufficiently small interaction in the iDEDM regime, as given by the conditions in Table 2. The densities of DM and DE evolve with the scale factor according to the expressions (2.14) and (2.15). This interaction was previously studied in [8, 16, 46, 134, 129, 171, 172, 131, 148, 38, 173, 132, 38, 174, 135, 175, 14, 176, 177, 139, 17, 178, 53, 55, 57, 136, 138, 179, 154, 155, 180, 16, 156, 161, 147, 181, 182, 162, 163, 164, 183, 166, 156, 147, 184, 185, 157, 158, 159, 160, 186, 187, 188, 189, 178, 190, 191, 150, 192, 193, 190, 194, 153, 195, 191, 196, 197, 198, 199, 200].
@@
$\begin{split}h(z)=\Bigg\{&\Omega_{\text{(dm,0)}}(1+z)^{3(1-\delta)}\\[2.84526pt] &+\Biggl(\Omega_{\text{(de,0)}}+\Omega_{\text{(dm,0)}}\left(\frac{\delta}{\delta+w}\right)\Bigl[1-(1+z)^{-3(\delta+w)}\Bigr]\Biggr)(1+z)^{3(w+1)}\\[2.84526pt] &+\Omega_{\text{(bm,0)}}(1+z)^{3}+\Omega_{\text{(r,0)}}(1+z)^{4}\Bigg\}^{\tfrac{1}{2}}.\end{split}$ (2.16)
@@ 2.1.5 Linear IDE model 5: $Q=3H\delta\rho_{\text{de}}$
This interaction mostly changes the dynamics for the late-time expansion and the distant future during DE domination. For this interaction, $Q=0$ if $\rho_{\text{de}}=0$; therefore, we can guarantee that $\rho_{\text{de}}\geq 0$ at all times. In contrast, this interaction will have future negative DM densities in the iDMDE regime, but this can be avoided with a sufficiently small interaction in the iDEDM regime, as given by the conditions in Table 2. The densities of DM and DE evolve with the scale factor according to the expressions (2.17) and (2.18). This interaction was previously studied in [134, 135, 16, 38, 129, 135, 201, 202, 131, 203, 204, 205, 148, 38, 132, 133, 38, 174, 135, 175, 14, 206, 207, 176, 177, 208, 203, 139, 209, 210, 19, 50, 178, 211, 21, 212, 213, 65, 55, 66, 51, 57, 214, 215, 58, 59, 67, 216, 62, 217, 136, 138, 154, 179, 155, 180, 16, 161, 218, 219, 181, 23, 182, 162, 163, 164, 199, 220, 208, 183, 156, 221, 58, 147, 184, 185, 157, 158, 159, 160, 186, 19, 222, 187, 189, 223, 211, 178, 190, 220, 62, 224, 191, 150, 225, 226, 153, 227, 196, 49, 228, 229, 230, 199, 231, 232, 68].
@@
$\begin{split}h(z)=\Bigg\{&\Biggl(\Omega_{\text{(dm,0)}}+\Omega_{\text{(de,0)}}\left(\frac{\delta}{\delta+w}\right)\Bigl[1-(1+z)^{3(\delta+w)}\Bigr]\Biggr)(1+z)^{3}\\[2.84526pt] &+\Omega_{\text{(de,0)}}(1+z)^{3(\delta+w+1)}+\Omega_{\text{(bm,0)}}(1+z)^{3}+\Omega_{\text{(r,0)}}(1+z)^{4}\Bigg\}^{\tfrac{1}{2}}.\end{split}$ (2.19)
@@ 2.2 Non-linear IDE models
$\begin{split}h(z)=\Bigg\{&\Bigl[\Omega_{\text{(dm,0)}}(1+z)^{3(1-\delta)}+\Omega_{\text{(de,0)}}(1+z)^{3(1+w)}\Bigr]\left[\frac{1+\left(\tfrac{\Omega_{\text{(dm,0)}}}{\Omega_{\text{(de,0)}}}\right)(1+z)^{-3(w+\delta)}}{1+\left(\tfrac{\Omega_{\text{(dm,0)}}}{\Omega_{\text{(de,0)}}}\right)}\right]^{-\tfrac{\delta}{w+\delta}}\\[2.84526pt] &+\Omega_{\text{(bm,0)}}(1+z)^{3}+\Omega_{\text{(r,0)}}(1+z)^{4}\Bigg\}^{\tfrac{1}{2}}.\end{split}$ (2.22)
@@
Similar to the kernel $Q=3H\delta\rho_{\text{dm}}$, this interaction primarily affects the dynamics in the distant past during DM domination but has a smaller impact on the late-time dynamics during DE domination. For this interaction, $Q=0$ if $\rho_{\text{dm}}=0$; therefore, we can guarantee that $\rho_{\text{dm}}\geq 0$ at all times. In contrast, this interaction leads to past negative DE densities in the iDMDE regime, which may be avoided with a small interaction in the iDEDM regime, as specified by the conditions in Table 2. The densities of DM and DE evolve with the scale factor according to the expressions (2.23) and (2.24). This interaction was previously studied in [233, 235, 16, 139, 140, 141, 173, 16, 161].
@@
$\begin{split}h(z)=\Bigg\{&\Biggl[\Omega_{\text{(dm,0)}}+\Omega_{\text{(de,0)}}\left(\frac{\Bigl[w+\delta\left(\tfrac{\Omega_{\text{(dm,0)}}}{\Omega_{\text{(de,0)}}}\right)\Bigr](1+z)^{3w}-\delta\left(\tfrac{\Omega_{\text{(dm,0)}}}{\Omega_{\text{(de,0)}}}\right)}{w}\right)\Biggr](1+z)^{3\left(1-\tfrac{w\delta}{w-\delta}\right)}\\[2.84526pt] &\times\left[\frac{\Bigl[w+\delta\left(\tfrac{\Omega_{\text{(dm,0)}}}{\Omega_{\text{(de,0)}}}\right)\Bigr](1+z)^{3w}+\left(\tfrac{\Omega_{\text{(dm,0)}}}{\Omega_{\text{(de,0)}}}\right)(w-\delta)}{w\Bigl[1+\tfrac{\Omega_{\text{(dm,0)}}}{\Omega_{\text{(de,0)}}}\Bigr]}\right]^{\tfrac{\delta}{w-\delta}}+\Omega_{\text{(bm,0)}}(1+z)^{3}+\Omega_{\text{(r,0)}}(1+z)^{4}\Bigg\}^{\tfrac{1}{2}}.\end{split}$ (2.25)
@@
Similar to the kernel $Q=3H\delta\rho_{\text{de}}$, this interaction primarily affects the dynamics of the late-time expansion and the distant future during DE domination, but has a smaller impact on the past dynamics during DM domination. For this interaction, $Q=0$ if $\rho_{\text{de}}=0$; therefore, we can guarantee that $\rho_{\text{de}}\geq 0$ at all times. In contrast, this interaction leads to future negative DM densities in the iDMDE regime, but this can be avoided with a sufficiently small interaction in the iDEDM regime, as specified by the conditions in Table 2. The densities of DM and DE evolve with the scale factor according to the expressions (2.26) and (2.27). This interaction was previously studied in [233, 235, 16, 139, 140, 141, 16, 161].
@@
$\begin{split}h(z)=\Bigg\{&\Biggl[\Omega_{\text{(dm,0)}}\left(\frac{\Bigl[w\left(\tfrac{\Omega_{\text{(dm,0)}}}{\Omega_{\text{(de,0)}}}\right)+\delta\Bigr](1+z)^{-3w}-\delta}{w\left(\tfrac{\Omega_{\text{(dm,0)}}}{\Omega_{\text{(de,0)}}}\right)}\right)+\Omega_{\text{(de,0)}}\Biggr](1+z)^{3\left(1+\tfrac{w^{2}}{w-\delta}\right)}\\[2.84526pt] &\times\left[\frac{\Bigl[w\left(\tfrac{\Omega_{\text{(dm,0)}}}{\Omega_{\text{(de,0)}}}\right)+\delta\Bigr](1+z)^{-3w}+w-\delta}{w\left(1+\tfrac{\Omega_{\text{(dm,0)}}}{\Omega_{\text{(de,0)}}}\right)}\right]^{\tfrac{\delta}{w-\delta}}+\Omega_{\text{(bm,0)}}(1+z)^{3}+\Omega_{\text{(r,0)}}(1+z)^{4}\Bigg\}^{\tfrac{1}{2}}.\end{split}$ (2.28)